OpenStudy (anonymous):

What is the exact value of the square root of the quantity 99 over 55.? Simplify if possible. 3 the square root of 5 all over 5. 3 the square root of 11 all over 5. 11 the square root of 3 all over 5. 5 the square root of 3 all over 5.

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

2x / (5x + 6) = 2 / x x = -1 and x = 6 both work 2(-1) / (5(-1) + 6) = 2 / (-1) - 2 / 1 = -2 / 1 2(6) / (5(6) + 6) = 2 / (6) 12 / 36 = 1 / 3 1/3 = 1/3